OpenStudy (anonymous):

Assume you have an annual income of $20,000 for each of the next 3 years and the inflation rate is 5% per year. Find the real value of your $20,000 salary for the next three years.

OpenStudy (anonymous):

I was looking through the book but couldn't find any exact way to solve it. However, the real way to solve this type of problem is if the CPI (Consumer Price Index) is given and there is an exact formula to adjust to inflation.

OpenStudy (anonymous):

I was wondering if I should take 5% of 20000 and add it on to each additional year. Please help.

OpenStudy (anonymous):

Year 1: Inflation 1.05 20000/1.05 = 19047.62 Year 2: Inflation 1.1025 20000/1.1025 = 18140.59 Year 3: Inflation 1.157625 20000/1.157625 = 17276.75 Add them up and you're done.

OpenStudy (anonymous):

Thank I sort of figured something like that before turning in the paper.